OpenStudy (zeroyang321):

Find the complex number z such that (1-4i)z is a real number and z+6-i is an imaginary number.

jhonyy9 (jhonyy9):

i think than use formula a^2 -b^2 =(a-b)(a+b) where a=1 and b= 4i will help you finding the z what will be equal (1+b) try it please and show your work

OpenStudy (jerry45):

@DanJS could you help with my question please

OpenStudy (danjs):

The same question is in Closed questions list from a bit ago http://openstudy.com/study#/updates/57fa098ee4b06a18dc414864

OpenStudy (kainui):

A number is real hen it's equal to its own complex conjugate. A number is imaginary when it's equal to the negative of its complex conjugate.

OpenStudy (phi):

I would let z= a+b i then compute (1-4i)(a+bi) = pure real and a+b i+6-i = pure imaginary notice in the 2nd equation we can write the sum as (a+6) + (b-1) i if it is pure imaginary, then the "real part" is zero. In other words a+6= 0 and we can "solve for a" now multiply out the first equation, and solve for the b value that makes the imaginary part of the product zero.

OpenStudy (kainui):

I might as well show my strategy, set up two equations based on \(real = real^*\) and \(imaginary = - imaginary^*\) : \[(1-4i)z=(1+4i)z^*\]\[(z+6-i)=-(z^*+6+i)\] second equation simplifies to: \[\frac{z+z^*}{2}=-6\]which is the real part of \(z\). first equation rearranges to: \[\frac{z-z^*}{2i} = 4\frac{z+z^*}{2}\]which says the imaginary part of z is 4 times the real part of z. Solved. :P